# Bridging Quantum Computing and Nuclear Structure: Atomic Nuclei on a Trapped-Ion Quantum Computer

### v1

@@ Abstract
We report accurate quantum simulations of medium-mass atomic nuclei—including oxygen, calcium, and nickel—on the RIKEN–Quantinuum Reimei trapped-ion quantum computer, achieving sub-percent accuracy. Using a symmetry-aware pair-unitary coupled-cluster doubles (pUCCD) ansatz implemented with a hard-core-boson mapping, and with particle-number–restoring post-selection, our ground-state energy estimates agree with noise-free statevector simulations and with exact diagonalization results on the order of $0.1\%$ relative error. These experiments establish a new benchmark for quantum simulations of strongly correlated nuclear systems on real hardware and demonstrate the near-term potential of high-fidelity trapped-ion platforms for nuclear-structure calculations.
Introduction
Quantum simulation of many-body systems is a leading application of quantum computing, with potential impact across quantum chemistry, condensed-matter physics and nuclear physics. While quantum algorithms and hardware have advanced rapidly, demonstrating clear advantage for interacting many-body problems remains an open challenge. For example, a recent study [1] identified scenarios in which quantum devices can outperform classical algorithms for certain spin models. Nonetheless, general claims of quantum advantage for many-body systems remain inconclusive and require careful, system-specific analysis.
On near-term devices, variational approaches such as the Variational Quantum Eigensolver [2, 3] rely crucially on the choice of an ansatz that balances implementability and expressivity. Common families of ansatze — including unitary coupled cluster (UCC) variants — can be effective in some domains but often suffer from optimization difficulties, called barren plateaus [4], or unfavorable scaling of parameters and gate depth as system size grows. This situation has evoked a broader consideration of quantum algorithms beyond NISQ towards fault-tolerant quantum computing (FTQC), leading to a term like Early Fault-Tolerant Quantum Computing (E-FTQC) [5].
In such situations, it is essential to broadly examine each component of quantum computation—namely, state preparation (also simply called the ansatz), the circuit for executing quantum algorithms, and the measurement strategy for evaluating the observable of interest—for each specific problem. The first part, the choice of the ansatz, plays a central role in any quantum algorithm. Even for fault-tolerant quantum computing (FTQC) approaches, such as Quantum Phase Estimation (QPE)[6] and early-FTQC algorithms [5], state preparation remains a fundamental building block for quantum algorithms. This is because the success of the quantum algorithm largely depends on the quality of the prepared quantum state, i.e., the overlap between the prepared state and the target state. In spin systems or chemical systems, various ansatze have been proposed, such as UCC-family, and ones based on matrix product states or tensor network states more generally [7, 8, 9].
Compared to other areas tackling quantum many-body problems, nuclear physics studies have seen relatively limited applications of quantum computing, despite several pioneering works [10, 11, 12], and recent reviews [13, 14]. In addition, only a few ansatze have been explored for nuclear systems, almost exclusively based on the unitary coupled cluster (UCC) family [15, 16, 12] with some variations such as Adapt-VQE [17]. A seminal work investigated the applicability of the UCC + Adapt-VQE to the nuclear shell model [12], where the number of CNOT gates required to achieve a given accuracy for shell-model calculations of oxygen and calcium isotopes was estimated.
Nuclear many-body systems are distinguished from other quantum many-body systems by exceptionally strong correlations arising from the nature of nuclear forces, posing a unique challenge for both classical and quantum computing. Even in the low-energy regime, where nucleons are the only explicit baryonic degrees of freedom, the problem is complicated by the fact that protons and neutrons—fermions of nearly equal mass—interact through multiple channels in an intricate manner. This results in a large number of non-negligible Hamiltonian terms and strong non-locality. Furthermore, anisotropic components of the interaction, such as tensor forces [18] and three-body forces [19], play an essential role in describing the rich phenomena of nuclear systems. These features are generally difficult to capture with a simple wave-function ansatz. Another manifestation of this complexity is entanglement entropy, which in nuclear systems obeys a volume law rather than an area law [20], hindering straightforward application of tensor networks, although extensions of tensor-network methods to such systems remain a promising direction [21, 22, 23].
Consequently, most quantum-computing studies of nuclear structure have focused on small systems or relied on classical simulators. Reported errors in ground-state energies range from $4\sim 13\%$ for 6Li [15] and $3\sim 13\%$ for oxygen isotopes [16]. In our previous work [24], we achieved relative errors below $0.1\%$ for 6He and 18O, and about $1\%$ for 42Ca on an IBM superconducting quantum device. However, those target systems were limited to valence two-neutron configurations and may not be scalable to heavier nuclei.
@@ Introduction
In this work, we introduce a hard-core-boson (HCB) mapping of the nuclear shell model combined with the pair-unitary coupled-cluster doubles (pUCCD) ansatz, which efficiently captures strong pairing correlations between nucleons of the same species. While pUCCD ansatze have been shown to be effective for chemical systems [25, 26, 27], this study reports, to our knowledge, the first demonstration of pUCCD applied to nuclear many-body systems on a quantum device. Implemented on the high-fidelity trapped-ion processor RIKEN–Quantinuum Reimei, and augmented by symmetry-aware ansatze and post-selection techniques, our approach yields ground-state energies for oxygen, calcium, and nickel isotopes that agree with noise-free statevector simulations within sub-percent relative error. These results establish a new experimental benchmark for quantum simulations of strongly correlated nuclear systems and demonstrate the near-term potential of trapped-ion platforms for nuclear-structure calculations.
Results
Figure 1: Ground-state energies of target isotopes on noise-free statevector simulators. a–c Full configuration interaction (FCI, dotted lines with gray circles) and doubly occupied configuration interaction (DOCI, blue diamonds) are compared with statevector simulations using the pUCCD (GS) ansatz (green stars) and the pUCCD (G) ansatz (red triangles). Here, pUCCD (GS) denotes the pUCCD ansatz implemented with nearest-neighbor Givens rotations and SWAP gates, while pUCCD (G) assumes all-to-all connectivity, as realized in the Reimei device. d–f Differences between the pUCCD ansatz results and the DOCI results are shown relative to FCI in logarithmic scale. Uncertainty bands for pUCCD (G) are obtained by random sampling over circuits with different initial configurations and Givens-rotation orderings.
In this section, we present the results of the pUCCD ansatz on the Reimei quantum hardware for the ground-state energies of oxygen, calcium, and nickel isotopes. The Reimei is based on Quantinuum’s H-1 series trapped-ion quantum computer, and has 20 qubits with all-to-all connectivity. The Reimei device is installed in RIKEN’s Wako campus, and started its operation in February 2025. Calculations are performed via Nexus [28], a cloud-based quantum computing platform provided by Quantinuum enabling users to run quantum simulations on various backends, including noise-free statevector simulators, emulators, and real devices.
Target systems
In this work, we study even-neutron isotopes of oxygen, calcium, and To demonstrate the capability of our approach, we focus on oxygen, calcium, and nickel isotopes, which span light to medium-mass nuclei and serve as standard benchmarks in nuclear structure calculations. The corresponding valence spaces are restricted to $sd$, pfpf, and $jj45$ shells, respectively, allowing feasible implementations on classical/quantum hardware and direct comparison with classical exact diagonalization.
As detailed in Methods section, we adopt the so-called pair unitary coupled cluster doubles (pUCCD) ansatz by folding the Fermionic degrees of freedom into Hard-core boson (HCB) degrees of freedom. The resulting qubit Hamiltonian takes the form:
$H_{\mathrm{qubit}}$ $=\sum_{i}\frac{\varepsilon_{i}+V_{ii}}{2}(I_{i}-Z_{i})+\frac{1}{4}\sum_{i\neq j}V_{ij}(X_{i}X_{j}+Y_{i}Y_{j})$
$+\frac{1}{4}\sum_{i\neq j}V^{m}_{ij}(I_{i}-Z_{i})(I_{j}-Z_{j}),$ (1)
All terms of the Hamiltonian Eq. (1) except the second are diagonal in the computational basis and can be measured directly by measuring the expectation values of the Pauli-$Z$ operators. For the non-diagonal $XX+YY$ terms, we employed two strategies. The first, called “Hadamard", uses Hadamard gates to measure $\langle XX\rangle$, from which $\langle YY\rangle$ follows by symmetry. The second, called “Basis rotations", applies a basis rotation to diagonalize $XX+YY$ in the computational basis; for this strategy, we additionally post-selected bitstrings to restore particle-number symmetry. More details on the measurement are provided in the Methods section.
Ideal simulation results of the pUCCD ansatz
We first assess the performance of the pUCCD ansatz on a noiseless statevector simulator, thereby probing its intrinsic accuracy. As a baseline, Full Configuration Interaction (FCI) provides the exact ground-state energies within the valence space. DOCI provides the exact solution within the zero-seniority space, whereas pUCCD offers a compact exponential parametrization of this space, and thus approximates DOCI. Therefore, any deviation from the FCI results reflects the intrinsic limitations of the ansatz itself, independent of hardware noise or statistical noise arising from the limited number of measurement shots.
In what follows, we consider two types of pUCCD ansatz, named pUCCD (GS) and pUCCD (G). The (GS) represents Givens plus SWAP, which consists of only nearest neighbor Givens rotations and thereby more suitable for hardware with a limited connectivity [29, 25]. On the other hand, (G) stands for Givens rotations where all-to-all connectivity is assumed. In experiments on the Reimei device, we will focus on the pUCCD (G) ansatz. For more details of ansatz design, see Methods section.
Fig. 1 shows ground-state energies for the target isotopes, obtained with pUCCD (G), pUCCD (GS), DOCI, and FCI on noise-free statevector simulators. At this scale, all curves are nearly indistinguishable, and the lower panels therefore highlight relative deviations from FCI on a logarithmic scale. Results for two-neutron (or two-neutron-hole) systems are omitted, since both DOCI and pUCCD exactly reproduce FCI in those cases.
For the pUCCD ansatze, the symbols show the results of one representative realization of the ansatz, and the uncertainty bands are plotted for the pUCCD (G) ansatz by random sampling of the circuit with different reference states and order of Givens rotations. The detailed explanation of the uncertainty bands is given in Supplementary Information.
Most importantly, pUCCD reproduces DOCI energies with an error of less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [30]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space (DOCI), whereas oxygen isotopes show larger contributions from higher-seniority configurations, leading to slightly larger deviations from FCI.
Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space, $jj45$ shell, consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [31]. The single-Trotter pUCCD ansatz, while efficient for capturing pairing correlations, becomes less expressive in the presence of such strong configuration mixing.
Results on RIKEN-Quantinuum’s Reimei device
Table 1: Summary of pUCCD results on the RIKEN–Quantinuum Reimei device. The “FCI” column lists exact ground-state energies obtained with the full configuration interaction method in the original fermionic representation. The “Ideal pUCCD” column shows results from the pUCCD ansatz in the hard-core-boson formalism. The fourth and fifth columns show experimental results on the Reimei device using two different post-selection techniques. The numbers in parentheses indicate the relative errors with respective to the pUCCD statevector simulations. The last column gives the valid ratio, i.e., the fraction of measured bitstrings preserving the correct particle number of the target nucleus. All energies are in MeV.
Nuclide FCI Ideal pUCCD Reimei: Hadamard Reimei: Basis rotation Valid ratio
18O -11.932 -11.932 -11.802 ($1.1\times 10^{0}$%) -11.893 ($3.3\times 10^{-1}$%) 0.98
20O -23.632 -22.939 -22.383 ($2.4\times 10^{0}$%) -22.854 ($3.7\times 10^{-1}$%) 0.98
22O -34.498 -33.924 -33.560 ($1.1\times 10^{0}$%) -33.829 ($2.8\times 10^{-1}$%) 0.97
24O -41.225 -41.021 -40.985 ($8.8\times 10^{-2}$%) -41.027 ($1.5\times 10^{-2}$%) 0.97
26O -40.869 -40.869 -40.613 ($6.3\times 10^{-1}$%) -40.940 ($1.8\times 10^{-1}$%) 0.96
42Ca -19.734 -19.734 -19.727 ($3.6\times 10^{-2}$%) -19.775 ($2.1\times 10^{-1}$%) 0.97
44Ca -38.675 -38.462 -38.540 ($2.0\times 10^{-1}$%) -38.424 ($9.9\times 10^{-2}$%) 0.97
46Ca -56.667 -56.440 -56.414 ($4.7\times 10^{-2}$%) -56.565 ($2.2\times 10^{-1}$%) 0.96
48Ca -73.662 -73.556 -73.538 ($2.5\times 10^{-2}$%) -73.489 ($9.1\times 10^{-2}$%) 0.95
50Ca -85.055 -84.921 -85.135 ($2.5\times 10^{-1}$%) -85.007 ($1.0\times 10^{-1}$%) 0.94
52Ca -95.360 -95.190 -95.038 ($1.6\times 10^{-1}$%) -95.149 ($4.3\times 10^{-2}$%) 0.94
54Ca -102.632 -102.557 -102.622 ($6.3\times 10^{-2}$%) -102.648 ($8.8\times 10^{-2}$%) 0.94
56Ca -104.589 -104.462 -104.620 ($1.5\times 10^{-1}$%) -104.325 ($1.3\times 10^{-1}$%) 0.94
58Ca -106.666 -106.666 -106.885 ($2.0\times 10^{-1}$%) -106.668 ($1.7\times 10^{-3}$%) 0.94
58Ni -21.447 -21.447 -21.207 ($1.1\times 10^{0}$%) -21.291 ($7.3\times 10^{-1}$%) 0.93
60Ni -41.276 -40.961 -40.752 ($5.1\times 10^{-1}$%) -40.992 ($7.4\times 10^{-2}$%) 0.94
62Ni -59.602 -59.013 -58.703 ($5.2\times 10^{-1}$%) -59.015 ($4.5\times 10^{-3}$%) 0.93
64Ni -76.444 -75.687 -76.027 ($4.5\times 10^{-1}$%) -75.765 ($1.0\times 10^{-1}$%) 0.92
66Ni -91.941 -91.059 -90.552 ($5.6\times 10^{-1}$%) -90.695 ($4.0\times 10^{-1}$%) 0.93
68Ni -106.140 -105.237 -105.169 ($6.5\times 10^{-2}$%) -105.288 ($4.8\times 10^{-2}$%) 0.92
70Ni -119.034 -118.215 -118.189 ($2.2\times 10^{-2}$%) -118.158 ($4.8\times 10^{-2}$%) 0.91
72Ni -130.930 -130.379 -130.511 ($1.0\times 10^{-1}$%) -130.397 ($1.3\times 10^{-2}$%) 0.95
74Ni -141.932 -141.694 -141.565 ($9.1\times 10^{-2}$%) -141.715 ($1.5\times 10^{-2}$%) 0.93
76Ni -152.103 -152.103 -151.986 ($7.7\times 10^{-2}$%) -152.054 ($3.2\times 10^{-2}$%) 0.94
Figure 2: Ground-state energies of target isotopes on the RIKEN–Quantinuum Reimei device. a oxygen, b calcium, c nickel isotopes. Energy estimates are shown relative to noise-free statevector simulations. The error bars represent statistical uncertainties, evaluated by bootstrapping with 500 resampling of datasets obtained from 1,024 measurement shots. The shaded band corresponds to a $0.3\%$ relative error with respect to the statevector results; most data points fall within this range.
We next implement the pUCCD (G) ansatz on the Reimei trapped-ion quantum computer. Table 1 summarizes the hardware results of the pUCCD (G) ansatz together with ideal simulations by noise-free statevector simulators. The measured energies agree closely with the noise-free simulations, showing relative errors mostly below $0.3\%$ and at most a few percent relative to FCI. These achievements represent the most accurate quantum-computing results for nuclear many-body ground states obtained on real devices to date.
Between the two measurement strategies, the basis-rotation method consistently outperforms the Hadamard method, yielding results closer to ideal simulations. This trend is further supported by bootstrapping analyses shown in Fig. 2. Lighter isotopes exhibit larger uncertainties because $XX+YY$ terms, where the difference between the two measurement strategies lies, contribute more significantly to the total energy in lighter isotopes. In contrast, for heavier isotopes, the mass-dependent scaling commonly applied to two-body matrix elements in phenomenological shell-model interactions reduces the absolute size of the XX+YY contributions (e.g., TBMEs for 58Ca and 78Ni are multiplied by $\sim 0.9$ compared with 42Ca and 58Ni), which further diminishes their relative contribution to the total energy.
Discussion
Quantum many-body systems has become a central focus of quantum computing research over the past decade, driven by rapid progress in algorithms and hardware. However, demonstrating genuine quantum advantage in interacting many-body physics remains an open challenge, requiring advances in all components of the algorithmic workflow: ansatz design, measurement strategies, and optimization techniques. In this work, we concentrated on ansatz construction and measurement from the standpoint of nuclear structure, where exceptionally strong correlations make the choice of ansatz especially critical.
By combining a hard-core boson (HCB) mapping of the nuclear shell-model Hamiltonian with a pair-unitary coupled cluster doubles (pUCCD) ansatz, we demonstrated accurate ground-state calculations of oxygen, calcium, and nickel isotopes on the RIKEN–Quantinuum Reimei trapped-ion quantum computer. Our results achieve sub-percent agreement with noise-free statevector simulations and exact diagonalization results, setting a new benchmark for nuclear-structure simulations on real NISQ hardware.
Although the pUCCD parameters were optimized using a classical simulator, this proof-of-principle study highlights how physically motivated, compact ansatze can deliver accurate results with modest qubit numbers and circuit depths. Such strategies may guide the discovery of efficient variational forms for larger systems where classical pre-optimization is infeasible.
Looking ahead, incorporating proton–neutron interactions will require ansatze capable of describing more general non-local correlations. Nonetheless, the present results demonstrate the viability of pUCCD for multi-neutron (or multi-proton) systems, and point toward applications to heavier systems such as neutron drops or neutron-rich isotopes. As an illustration of possible extensions, entropy-driven entanglement forging [32] has recently been proposed to localize correlations in Adapt-VQE circuits; similar ideas could be combined with pUCCD, for instance by monitoring generalized seniority.
Beyond ansatz design, new algorithmic paradigms merit exploration. Quantum subspace diagonalization techniques [33, 34, 35, 36, 37], including Quantum Krylov and related methods, use quantum devices to generate subspace states while delegating diagonalization to classical post-processing. These approaches provide access to excited states and transition properties such as electromagnetic transitions. Such hybrid methods, combined with compact ansatze like pUCCD, represent a promising path toward realistic quantum simulations of nuclear spectroscopy in the near future.
Methods
Basis states and Hamiltonians
@@ Methods
or systems within one or two major shells, as in this work, the one-body term reduces to the diagonal form $\varepsilon_{i}\equiv h_{ii}$, called the single-particle energies (SPEs). The fermionic creation and annihilation operators, $a^{\dagger}_{i}$ and $a_{i}$, act on these single-particle states, while $V_{ijkl}$ denotes the two-body matrix elements (TBMEs).
These SPEs and TBMEs can be either derived from the realistic nucleon-nucleon interactions or from the phenomenological adjustments to reproduce experimental data for binding energies, excitation energies, and so on. The model space and adopted interactions in this work are summarized in Table 2. The methods in this work can be applied to any other interactions including microscopically derived ones. Diagonalization with these interactions can be reproduced using publicly available codes such as KSHELL [38, 39] or NuclearToolkit.jl [40, 41].
Table 2: Adopted interactions and model spaces.
@@
Once the Hamiltonian is defined, one can diagonalize the Hamiltonian in a given model space, leading to the Configuration Interaction (CI) method. Here, "full configuration interaction" (FCI) refers to exact diagonalization within the selected valence space, following the convention of quantum chemistry (rather than the no-core shell model definition often used in nuclear physics).
@@
It should also noted that the way to represent the Hamiltonian in paired form is not unique, and the choice of the pairs can be different depending on the context. In the case of electronic structure calculations, the natural assumption is to consider the spin up and spin down electrons as a pair. Detailed derivations of the zero-seniority Hamiltonian for electronic structure calculations can be found in e.g. [46]. For nuclear many-body methods one has more freedom involved in the system, and one can consider the pairs of nucleons with different spins and/or different orbitals.
Fig. 3 shows the schematic figure of the hard-core boson (HCB) representation, which is a core assumption in this work for describing target systems efficiently. As ab example, the figure shows a configuration of 22O nucleus on top of the inert 16O core. The single-particle states are specified by the quantum numbers $\{n,l,j,j_{z},t_{z}\}$, leading to 12 single-particle states for both protons and neutrons in the $sd$ shell.
Within the zero-seniority approximation, the active degrees of freedom can be restricted to time-reveral pairs of nucleons (neutrons in this case). Consequently, the number of qubits required to represent the system is reduced to half the number of single-particle states in the original fermionic representation.
In the HCB representation, the Hamiltonian becomes
@@
where $i<0$ denotes single-particle states with negative $j_{z}$ values, and $\bar{i}$ is the time-reversal partner of the single-particle state $i$. The last term, absent in the valence two-neutron systems studied previously [24], can be regarded as a mean-field conttibution from other occupied pairs.
Defining pair creation, annihilation, and occupation operators,
@@
the Hamiltonian can be rewritten in pairwise or hard-core boson (HCB) form as
$H^{\mathrm{pw}}$ $=\sum_{\tilde{i}}(\varepsilon_{\tilde{i}}+V^{p}_{\tilde{i}\tilde{i}})N_{\tilde{i}}+\sum_{\tilde{i}\neq\tilde{j}}V^{p}_{\tilde{i}\tilde{j}}A^{\dagger}_{\tilde{i}}A_{\tilde{j}}+\sum_{\tilde{i}\neq\tilde{j}}V^{m}_{\tilde{i}\tilde{j}}N_{\tilde{i}}N_{\tilde{j}}$, (7)
@@
$V^{m}_{\tilde{i}\tilde{j}}$ $\equiv V_{ijij}+V_{i\bar{j}i\bar{j}}+V_{\bar{i}j\bar{i}j}+V_{\bar{i}\bar{j}\bar{i}\bar{j}}.$ (10)
@@
More details on this derivation are provided in the Supplementary Information. The factor $1/2$ in Eq. (6) is sometimes omitted (as in our earlier work [24]), but is retained here to emphasize the intererpretaion of $N_{\tilde{i}}$ as the occupation number of a pair degree of freedom.
The paired-neutron operators satisfy bosonic commutation relations, as detailed in e.g. [46, 25, 24], but multiple occpancy of the same pair is forbidden, hence the term hard-core boson (HCB). A key advantage of the HCB representation—beyond the qubit reduction—is the elimination of non-local Jordan–Wigner strings when mapping to Pauli operators. From Jordan-Wigner mappings, Pauli-$Z$ terms appear to maintain the anti-commutation relations of the fermionic operators. In some literature, these Pauli-$Z$ terms are omitted, which is referred to as qubit excitation in Refs. [48, 49], as a hardware-efficient mapping. Those Pauli-$Z$ terms naturally cancel out in the pair-wise form of the Hamiltonian. Hence, one can map the Hamiltonian to the Pauli operators into simple forms as Eq. (1).
All the terms fall into two categories: One is diagonal terms in the computational basis, $\{I,Z,ZZ\}$, measurable directly, and the other is off-diagonal terms, $\{XX,YY\}$, requiring additional measurement circuits. The measurement strategies are described later in this section.
ansatze
The ansatz plays a central role in quantum algorithms, both in the NISQ and FTQC regimes. In practice, the ansatz (i.e., the trial wave function) must satisfy two requirements: (1) it should be implementable on available quantum hardware, and (2) it should be expressive enough to approximate, e.g., the true ground state with high fidelity.
In this work, we employ the pair unitary coupled cluster doubles (pUCCD) ansatz for even-even multi-neutron systems. As shown in Fig. 1, pUCCD provides ground-state energies comparable in accuracy to doubly occupied configuration interaction (DOCI) method, while requiring far fewer parameters.
The pUCCD ansatz is a quantum version of the pair coupled cluster doubles method. A seminal work on the pUCCD ansatz in the context of quantum computing is found in Ref. [25], where strategies for measurement and post-selection were also proposed. In chemistry, pUCCD has been successfully applied to systems of paired electrons with spin up and spin down [25, 26, 27].

### v2

@@ Abstract
We demonstrate quantum simulations of strongly correlated nuclear many-body systems on the RIKEN-Quantinuum Reimei trapped-ion quantum computer, targeting ground states of oxygen, calcium, and nickel isotopes. By combining a hard-core-boson representation of the nuclear shell model with a pair-unitary coupled-cluster doubles ansatz, we achieve sub-percent relative error in the ground-state energies compared to noise-free statevector simulations. Our approach leverages symmetry-aware state preparation and particle-number post-selection to efficiently capture pairing correlations characteristic of systems with same-species nucleons. These findings highlight the viability of high-fidelity trapped-ion platforms for nuclear physics applications and provide a foundation for scaling to more complex nuclear systems.
I Introduction
Quantum simulation of many-body systems is a leading application of quantum computing, with potential impact across quantum chemistry, condensed-matter physics, nuclear physics and so on. While quantum algorithms and hardware have advanced rapidly, demonstrating clear advantage for interacting many-body problems remains an open challenge. For example, a recent study [1] identified scenarios in which quantum devices can outperform classical algorithms for certain spin models. Nonetheless, general claims of quantum advantage for many-body systems remain inconclusive and require careful, system-specific analysis.
On near-term devices, variational approaches such as the Variational Quantum Eigensolver [2, 3] rely crucially on the choice of an ansatz that balances implementability and expressivity. Common families of ansatze — including unitary coupled cluster (UCC) variants — can be effective in some domains but often suffer from optimization difficulties, called barren plateaus [4], or unfavorable scaling of parameters and gate depth as system size grows. This situation has evoked a broader consideration of quantum algorithms beyond NISQ towards fault-tolerant quantum computing (FTQC), leading to a term like Early Fault-Tolerant Quantum Computing (E-FTQC) [5]. In such situations, it is essential to broadly examine each component of quantum computation—namely, state preparation (also simply called the ansatz), quantum algorithms, and measurement strategies for evaluating the observable of interest—for each specific problem.
The first part, the choice of the ansatz, is particularly crucial. This is because the success of most of quantum algorithms for quantum many-body problems largely depends on the quality of the prepared quantum state, i.e., the overlap between the prepared state and the target state. In spin systems or chemical systems, various ansatze have been proposed, such as UCC-family, and ones based on matrix product states or tensor network states more generally [6, 7, 8]. Compared to other areas tackling quantum many-body problems, nuclear physics studies have seen relatively limited applications of quantum computing, despite several pioneering works [9, 10, 11], and recent reviews [12, 13]. To date, only a few ansatze have been explored for nuclear systems, almost exclusively based on the unitary coupled cluster (UCC) family [14, 15, 11] with some variations such as Adapt-VQE [16]. A seminal work investigated the applicability of the UCC + Adapt-VQE to the nuclear shell model [11], where the number of CNOT gates required to achieve a given accuracy for shell-model calculations of oxygen and calcium isotopes was estimated.
Nuclear many-body systems are distinguished from other quantum many-body systems by exceptionally strong correlations arising from the nature of nuclear forces, posing a unique challenge for both classical and quantum computing. Even in the low-energy regime, where nucleons are the only explicit baryonic degrees of freedom, the problem is complicated by the fact that protons and neutrons—fermions of nearly equal mass—interact through multiple channels in an intricate manner. This results in a large number of non-negligible Hamiltonian terms and strong non-locality. Furthermore, anisotropic components of the interaction, such as tensor forces [17] and three-body forces [18], play an essential role in describing the rich phenomena of nuclear systems. These features are generally difficult to capture with a simple wave-function ansatz. Another manifestation of the complexity is entanglement entropy, which in nuclear systems obeys a volume law rather than an area law [19], hindering straightforward application of tensor networks.
Consequently, most quantum-computing studies of nuclear structure have focused on small systems or relied on classical simulators. Reported errors in ground-state energy estimates on quantum hardware range from $4-13\%$ for 6Li [14] and $3-13\%$ for oxygen isotopes [15]. In our previous work [20], we achieved relative errors below $0.1\%$ for 6He and 18O, and about $1\%$ for 42Ca on an IBM superconducting quantum device. However, the target systems were limited to valence two-neutron configurations and our method is not straightforwardly extendable to more general nuclear systems. This motivates the present work to develop more general and scalable approaches.
In this work, we introduce a hard-core-boson (HCB) mapping of the nuclear shell model combined with the pair-unitary coupled-cluster doubles (pUCCD) ansatz, which efficiently captures strong pairing correlations between nucleons of the same species. While pUCCD ansatze have been shown to be effective for chemical systems [21, 22, 23], this study reports, to our knowledge, the first demonstration of pUCCD applied to nuclear many-body systems on a quantum device. Implemented on the high-fidelity trapped-ion processor RIKEN-Quantinuum Reimei, and augmented by symmetry-aware ansatze and post-selection techniques, our approach yields ground-state energies for oxygen, calcium, and nickel isotopes that agree with noise-free statevector simulations within sub-percent relative error. These results establish a new experimental benchmark for quantum simulations of strongly correlated nuclear systems and demonstrate the near-term potential of trapped-ion platforms for nuclear-structure calculations.
This paper is organized as follows. In Sec. II, we describe the methods used in this work, including the Hamiltonian mapping, the adopted ansatz, and the measurement strategies. In Sec. III, we present the results for oxygen, calcium, and nickel isotopes, noise-free simulations and hardware experiments. Finally, we summarize our findings and discuss future outlooks in Sec. IV.
II Methods
II.1 Target systems and Hamiltonian mapping
@@
For systems within single major shell, as in this work, the one-body term reduces to the diagonal form $\varepsilon_{i}\equiv h_{ii}$, called the single-particle energies (SPEs). The fermionic creation and annihilation operators, $a^{\dagger}_{i}$ and $a_{i}$, act on these single-particle states, while $V_{ijkl}$ denotes the two-body matrix elements (TBMEs).
These SPEs and TBMEs can be either derived from the realistic nucleon-nucleon interactions or from the phenomenological adjustments to reproduce experimental data for binding energies, excitation energies, and so on. The target nuclei, model space and adopted interactions in this work are summarized in Table 1. The methods in this work can be applied to any other interactions including microscopically derived ones as far as one considers similar systems with same-species nucleons under the methods described below. Diagonalization with these interactions can be reproduced using publicly available codes such as KSHELL [24, *KSHELL2] or NuclearToolkit.jl [26, *Repo_NuclearToolkit.jl].
Table 1: Target nuclei, interactions, and model spaces.
@@
Once the Hamiltonian is defined, one can diagonalize the Hamiltonian in a given model space by means of variants of Configuration Interaction (CI) methods, known as the nuclear shell model in nuclear physics community. In this work, full configuration interaction (FCI) refers to exact diagonalization within the selected valence space, following the convention of quantum chemistry (rather than the no-core shell model definition often used in nuclear physics).
@@
In the case of electronic structure calculations, the natural assumption on the pair degree of freedom is to consider the spin up and spin down electrons as a pair. Detailed derivations of such zero-seniority Hamiltonian for electronic structure calculations can be found in e.g. [33]. For nuclear many-body methods one has more freedom involved in the system, so one can consider the pairs of nucleons with different spins and/or different orbitals. For this reason, we will detail the way to introduce the paired form of a nuclear shell-model Hamiltonian.
The hard-core-boson (HCB) representation provides a natural way to introduce paired degrees of freedom by folding each time-reversed single-particle pair $(i,\bar{i})$ into a single boson-like site that can be occupied or empty but forbids multiple occupancy. Figure 1 illustrates this mapping for 22O on top of the 16O inert core: the valence single-particle states are grouped into pair orbitals, which reduces the number of active degrees of freedom (and hence qubits). The single-particle states in the original fermionic representation are specified by the quantum numbers $\{n,l,j,j_{z},t_{z}\}$, leading to 12 single-particle states for both protons and neutrons in the $sd$ shell. In the HCB representation, these 12 single-particle states form 6 pair orbitals for neutrons, each represented by a qubit. Focusing on even-number neutron systems and terms involving time-reversed pairs, the shell-model Hamiltonian can be expressed as
@@
where $i<0$ denotes single-particle states with negative $j_{z}$ values, and $\bar{i}$ is the time-reversal partner of the single-particle state $i$. The last term, absent in the valence two-neutron systems studied previously [20], can be regarded as a mean-field contribution from other occupied pairs. Then, the Hamiltonian can be expressed in terms of the paired operators defined below:
@@
where $A^{{\dagger}}_{\tilde{i}}$ and $A_{\tilde{i}}$ are the pair creation and annihilation operators, and $N_{\tilde{i}}$ is the pair occupation number operator, $\tilde{i}$ denotes the pair index corresponding to the single-particle states $i$ and $\bar{i}$. Then, the Hamiltonian in Eq. (2) can be rewritten as
$H_{\mathrm{HCB}}=$ $\sum_{\tilde{i}}(\varepsilon_{\tilde{i}}+V^{p}_{\tilde{i}\tilde{i}})N_{\tilde{i}}$
$+\sum_{\tilde{i}\neq\tilde{j}}V^{p}_{\tilde{i}\tilde{j}}A^{\dagger}_{\tilde{i}}A_{\tilde{j}}+\sum_{\tilde{i}\neq\tilde{j}}V^{m}_{\tilde{i}\tilde{j}}N_{\tilde{i}}N_{\tilde{j}}$, (6)
@@
$V^{m}_{\tilde{i}\tilde{j}}\equiv$ $V_{ijij}+V_{i\bar{j}i\bar{j}}+V_{\bar{i}j\bar{i}j}+V_{\bar{i}\bar{j}\bar{i}\bar{j}},$ (9)
@@
where $V^{p}$ and $V^{m}$ are introduced to emphasize the pairing and monopole nature of the interactions, respectively. The factor $1/2$ in Eq. (5) is sometimes omitted (as in our earlier work [20]), but is retained here to interpret $N_{\tilde{i}}$ as the occupation number of a pair degree of freedom. More details on this derivation are provided in the Supplemental Material [35].